find the derivative of 2(pi)(r)-(2v/r^2): find the derivative of 2(pi)(r)-(2v/r^2)
Answer by bucky(1732)

2(pi)(r)-(2v/r^2)
.
The problem asks for you to take the derivative of the above expression. I presume that
the derivative is with respect to r and that v is a constant.
.
One term at a time. The product 2(pi) is a constant so it is a multiplier of the derivative
of r. The rule that applies is that the derivative of r^n is n*r^(n-1)
In this first term the exponent of r is 1. So you are taking the derivative of r^1 which
the rule tells you in 1*r^(1-1) and this simplifies to 1*r^0 = 1*1 = 1.
Putting this all together for the first term, the derivative is 2*(pi)*1 = 2*(pi).
.
On to the second term. The most critical point here is to recognize that a positive
exponent in the denominator is equivalent to a negative exponent in the numerator.
Using this we can convert (2v)/r^2 to an equivalent form (2v)*r^(-2).
.
Now we can apply the same technique as we did for the first term. The factor (2v) is
presumed to be a constant and therefore will be a multiplier of the derivative of r^(-2).
Again we use the rule that the derivative of r^n is n*r^(n-1).
So the
derivative of r^(-2) is (-2)*r^(-2-1) which simplifies to (-2)*r^(-3).
Don't forget that this gets multiplied by the constant (2v) so that the derivative
for this second term is (2v)(-2)r^(-3). Multiplying this out results in:
.
(-4vr^(-3)) and since a negative exponent in the numerator becomes a positive
exponent in the denominator, we could also write the derivative as:
.
-4v/(r^3).
.
Finally we combine the derivatives for the first and second terms to get:
.
2(pi) - (-4v)/r^3
.
and taking care of the signs for the second term concludes the effort by producing
the result:
.
2(pi) + (4v)/r^3
.
Hope this helps you to see your way through the problem.